Superbowl contender: The probability that San Francisco plays in the next SuperBowl is nine times teh probability that they do not play in the Super Bowl. The probability that San Francisco plays in the next Super Bowl plus the probability that they do not play is 1. What is the probablility that San Francisco plays in the next Super Bowl?
I use the P=9 and the P=9w I got the answer fractioned out 1 out 9 they have a chance however how can this be place as a fraction problem to get a percent if I take 9divided by 9 i get 1.

Answer by stanbon(75887)

Superbowl contender: The probability that San Francisco plays in the next SuperBowl is nine times the probability that they do not play in the Super Bowl. The probability that San Francisco plays in the next Super Bowl plus the probability that they do not play is 1. What is the probablility that San Francisco plays in the next Super Bowl?
I use the P=9 and the P=9w I got the answer fractioned out 1 out 9 they have a chance however how can this be place as a fraction problem to get a percent if I take 9divided by 9 i get 1.
-------------------------------
Let P(not plays) = x ; Then P(play) = 9x
----
Equation x + 9x = 1
10x = 1
x = 1/10 = 10/100 = 10%
====================================
Cheers,
Stan H.